a communication signal is given by the function y=sint/t
derive and equation for dy/dt using the quotient rule.
1
Expert's answer
2020-04-08T03:29:42-0400

Quotient rule:

"d\left(\frac{u}{v}\right)=\frac{v\cdot du-u\cdot dv}{v^2}"


Given function

"y=\frac{sint}{t}"

Differentiating with respect to t using quotient rule

"\frac{dy}{dt}=\frac{t\cdot \frac{d}{dt}\left(sint\right)-sint\cdot \frac{d}{dt}\left(t\right)}{t^2}"

we know "\frac{d}{dt}\left(sint\right)=cost\:\:and\:\frac{d}{dt}\left(t\right)=1"

"\frac{dy}{dt}=\frac{t\cdot cost-sint\cdot 1}{t^2}"

"\frac{dy}{dt}=\frac{t\cdot cost-sint}{t^2}"